When 0.187 g of toluene, C7H8, is burned in a bomb calorimeter, the temperature of both the water and the calorimeter rises by 4.83 ∘C.

Assuming that the bath contains 250.0 g of water and that the heat capacity for the calorimeter is 525 J/∘C, calculate the combustion energy (ΔE) for toluene in kilojoules per gram.
Answer #1

When 0.187 g of toluene, C7H8, is burned in a bomb calorimeter, the temperature of both the water and the calorimeter rises by 4.83 ?C.

Assuming that the bath contains 250.0 g of water and that the heat capacity for the calorimeter is 525 J/?C, calculate the combustion energy (?E) for toluene in kilojoules per gram.

dE = -Qrxn / mass

dT = 4.83°C

Q = water + calorimeter = mwater*Cwater*dT + Ccal*dT

Q = 250*4.184*4.83 + 525*4.83 = 7587.93 J

this is lost by the system so

Qrxn = -Q = -7587.93 J

so

dE = Qrxn/m = 7587.93 / 0.187

dE = 40577.16 J/g

dE = 40.57 kJ/g